OpenStudy (anonymous):

Hard Problem** Double Integral volume of tetrahedron bounded by 4x+4y+7z=112 x=15y x=0 z=0

OpenStudy (anonymous):

image included

OpenStudy (anonymous):

k so i have the plane's intercepts at y = -16, x=-26, and z = -16

OpenStudy (anonymous):

not exactly sure how to set up the limits of integration

OpenStudy (anonymous):

so wait it would have to be split up with two integrals then

OpenStudy (nottim):

Hey, I'm thnking layman's method here, so don't trust me. Er..have you tried just pluggin in the values?

OpenStudy (nottim):

aww wait. I see how complex this is. So, vectors is not my thing.

OpenStudy (anonymous):

\[\int\limits_{-26}^{-1680/67}\int\limits_{-4/7*x-16}^{0}(-4/7)*x-y-16dydx\]

OpenStudy (anonymous):

that should be the first equation

OpenStudy (anonymous):

then add + \[\int\limits_{-1680/67}^{0}\int\limits_{(1/15)x}^{0}(-4/7)x-y-16dydx\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

i got this by finding the equation of the line that goes from the intercepts y=-16 to x=-26 **basically just the equation of the plane 4x+7y+7z=112 and setting z=0

OpenStudy (anonymous):

then the problem is that it intercepts y=1/15x so you'd have to find the interception of above with y=1/15x aka 1/15x=-4/7x-16 again, where -4/7x-16 is from taking the above equation and setting z = 0 and solving for y

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

answer = -(1225024/9849)

OpenStudy (anonymous):

could very easily be completely wrong on this

OpenStudy (nottim):

Sam, please say you got something on this.

OpenStudy (anonymous):

:0

OpenStudy (anonymous):

it can't have negative area tho right?

OpenStudy (anonymous):

er wait it should because it's in the um 8th quadrant

OpenStudy (anonymous):

er crap yeah just put it in and it's wrong

OpenStudy (nottim):

Honestly, I have no idea.

OpenStudy (anonymous):

yeah, think i'm going to have to take this to the math lab guarantee everyone is going to get this one wrong

OpenStudy (anonymous):

eh that's ok, thank you for the help

OpenStudy (nottim):

yeah. sorry bro

OpenStudy (calculator):

u can try ask yahoo answers